"-400,000 + 100,000/IRR - 3% = 0 = IRR =?"

I know the correct answer is 28%, I just don't remember how to solve this equation. Thank you!
Answer #1

Answer:

There is a bracket missing. (If the bracket is not there then answer will not be 28%)

It should be:

-400,000 + 100,000/ (IRR - 3%) = 0

Given:

-400,000 + 100,000/ (IRR - 3%) = 0

=> 100,000 / (IRR - 3%) = 400,000

=> 400,000 * (IRR - 3%) = 100,000

=> IRR - 3% = 100,000 / 400,000

=> IRR - 3% = 25%

=> IRR = 25% + 3%

=> IRR = 28%